OpenStudy (anonymous):

If log base a 5=x and log base a 7=y, then log base a sqrt of 1.4=?

OpenStudy (anonymous):

4 is your answer.

OpenStudy (irishboy123):

play with these, especially 2 & 3 

OpenStudy (irishboy123):

and note that \(1.4 = \dfrac{7}{5}\) so \(\sqrt{1.4} = 1.4^{\frac{1}{2}} = (\dfrac{7}{5})^{\frac{1}{2}}\) that's all you need

OpenStudy (anonymous):

oh that makes sense now. thanks!

OpenStudy (irishboy123):

cool :p